OpenStudy (anonymous):

2P2+3P2+4P2+5P2+....+24P2+25P2

OpenStudy (anonymous):

\[(\frac{n(n+1)}{2}-1)p2\]

Parth (parthkohli):

\[\large \sum_{i = 2}^{25}P(i,2)\]This is a good question.

OpenStudy (eyust707):

n = 25

Parth (parthkohli):

Careful, P means "permutations".

OpenStudy (anonymous):

Ya never can tell.. sometimes these guys just mean squared.... but I bet you're right.

OpenStudy (anonymous):

If @ParthKohli is correct... I entered sum(seq( x nPr 2 , x ,2 ,25)) into my TI-84 and got an answer.

hartnn (hartnn):

m getting \[\sum_{2}^{25}n(n-1)\]

OpenStudy (anonymous):

\[1\times2+2\times3+...+24\times25\]

hartnn (hartnn):

or \[\sum_{1}^{24}n(n+1)\]

hartnn (hartnn):

this can be solved.

OpenStudy (anonymous):

@hartnn I got the same answer with your minus version

hartnn (hartnn):

\[\sum_{1}^{24}n^2+n\]

OpenStudy (anonymous):

\[=\frac{24(24+1)(48+1)}{6}+\frac{24.25}{2}\]

hartnn (hartnn):

yup, thats it ^^

OpenStudy (anonymous):

hah, and plus version

hartnn (hartnn):

4*25*49+12*25=5200 did u get 5200 @EulerGroupie ?

OpenStudy (anonymous):

I did a sum of sequences of permutations in my calculator and got 5200

hartnn (hartnn):

@diazmayendra is not even here to understand this.

OpenStudy (anonymous):

@diazmayendra medal'd me and moved on to the next question

hartnn (hartnn):

okk....

OpenStudy (anonymous):

but I like it when people back these things up with more info... that's when I learn on stuff that I am weak on.

OpenStudy (anonymous):

new site added to favorites, thank you @ParthKohli

Parth (parthkohli):

@EulerGroupie the Mathematics on that site is advanced :)

OpenStudy (anonymous):

i not understand this question and i not understand too your answer